# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v2

@@ Abstract
Methods. We trained a deep learning (DL) based method Zoobot to estimate the fractional contribution of a central point source to the total light. Our training sample comprises realistic mock JWST images of simulated galaxies from the IllustrisTNG cosmological hydrodynamical simulations. We injected different amounts of the observed JWST point spread functions to represent galaxies with varying levels of AGN contribution. Galaxies in our training sample span a wide range of morphologies, including mergers. We analyse in detail the performance of our method as a function of various galaxy properties and compare with results obtained from the traditional light profile fitting tool GALFIT. After training, we applied our method to real JWST observations.
Results. We find excellent performance of our DL method in recovering the injected $f_{AGN}$, in terms of precision and accuracy. The mean difference between the predicted and true $f_{AGN}$ is $-0.002$ and the overall root mean square error (RMSE) is $0.013$. The overall relative absolute error (RAE) is $0.076$ and the outlier (defined as predictions with RAE $>20\%$) fraction is $6.5\%$. In comparison, using GALFIT, we achieve a mean difference of -$0.02$, RMSE of $0.12$, RAE of $0.19$ and outlier fraction of $19\%$. We also investigate how these key performance metrics obtained from Zoobot and GALFIT vary as a function of the injected $f_{AGN}$, redshift, signal-to-noise ratio, and galaxy size. In addition to the superior performance, our DL method has several other advantages over traditional methods. For example, it has a much higher success rate (even for highly disturbed or irregular galaxies) and is extremely fast. We applied our trained DL model to real JWST observations and found that 20% of the X-ray-selected AGN and 8% of the MIR-selected AGN are also identified as AGN using a cut at $f_{\rm AGN}>0.2$. When using $f_{\rm AGN}>0.1$, these overlaps increase to 33% for the X-ray AGN and 15% for the MIR AGN. In summary, our DL-based method to identify AGN and estimate AGN contribution fraction has a huge potential in future applications to large galaxy imaging surveys.
@@ 1 Introduction
It is crucial to correctly separate the central AGN light from the host galaxy, across a wide range of galaxy types and redshifts. Good quality optical imaging, with high spatial resolution and signal-to-noise ratio (S/N), is needed to decompose the observed total light into contributions from the host galaxy and the central AGN. Traditionally, this is done in photometric data by fitting two-dimensional (2D) profiles to the galaxy’s light, using one (or more if needed) analytic profile to describe the galaxy (typically a Sérsic profile) and a point spread function (PSF) profile to describe the central point source. Many studies use GALFIT (Peng et al., 2002), one of the most widely used 2D surface brightness modelling software to perform image-based decomposition of AGN and host galaxy light (e.g. Kim et al., 2008; Bentz et al., 2009; Gabor et al., 2009; Böhm et al., 2013; Schramm & Silverman, 2013; Du et al., 2014; Urbano-Mayorgas et al., 2019; Son et al., 2022; Aird et al., 2022; Ji et al., 2022; Dewsnap et al., 2023; Zhuang & Ho, 2023; Sturm & Reines, 2024; Zhuang et al., 2024). However, there are some known technical issues with GALFIT. For example, in certain circumstances, particularly for extremely compact sources, the minimisation algorithm used can be trapped in a local minimum. Other software have been used in an attempt to improve on GALFIT performance, such as PSFMC (Mechtley, 2014), which is a Markov Chain Monte Carlo (MCMC) simultaneous fitting software to perform multi-component profile fitting (Mechtley et al., 2016; Marshall et al., 2021) and LENSTRONOMY (Birrer et al., 2015; Birrer & Amara, 2018), which uses particle swarm optimisation (Kennedy & Eberhart, 1995) for $\chi^{2}$ minimisation, to reduce the likelihood of getting trapped in a local minimum when searching the parameter space, and MCMC for Bayesian parameter inference (Foreman-Mackey et al., 2013). The latter method is used by (Li et al., 2021) to decompose a sample of X-ray-selected AGN into quasars and host galaxy components, to investigate the properties of the host galaxies. However, these approaches all assume that the galaxy’s surface brightness profile can be well fitted by a single Sérsic profile (or combination of Sérsic profiles), which may not be the case, particularly for irregular galaxies, highly disturbed merging galaxies, or galaxies with complicated substructures.
@@
Consequently, if Sérsic profiles cannot adequately describe the host galaxy light, it can introduce a systematic bias in the derived luminosity of the PSF component and of the host galaxy. In addition, if there is a significant contribution to the total flux from the central AGN, it can lead to a bias in the morphology of the host galaxy, as it may appear more bulge-dominated (Pierce et al., 2010). Indeed, it is not always a good idea to represent a galaxy with a single Sérsic profile. Bentz et al. (2009) study AGN host galaxies at redshift $z\approx 0.7$, by fitting their surface brightness distributions with a combination of Sérsic profiles (to account for the host galaxy’s light) and PSF profile (to describe the AGN component). They find that most galaxies in their sample are well described with two Sérsic profiles, in addition to the PSF profile, and a few of them required three or more Sérsic profiles to describe additional components such as bars. Another possible difficulty in decomposing the AGN from the host galaxy is the (sometimes significant) variation of the PSF in a given survey due to spatial/temporal changes or differences in galaxy spectral energy distributions (SEDs), as an incorrect PSF can bias the estimated contribution of the AGN to the total flux of the galaxy. Kim et al. (2008) perform 2D decomposition of Sérsic + PSF profiles to galaxies hosting AGN in Hubble Space Telescope (HST) images and investigates the effect of realistic PSF mismatch, finding a systematic overestimation of the flux of the host galaxies, particularly for those containing bright AGN.
@@
Throughout the paper we assume a flat $\Lambda$CDM universe with $\Omega_{M}=0.2865$, $\Omega_{\Lambda}=0.7135$, and $H_{0}=69.32$ km s-1 Mpc-1 (Hinshaw et al., 2013). Unless otherwise stated, magnitudes are presented in the AB system.
@@ 2.1 JWST/COSMOS-WEB
Figure 1: Overview of the JWST/NIRCAM F150W PSFs in COSMOS-Web. We stacked all available PSFs from Zhuang et al. (2024) and show the mean PSF (left), standard deviation (centre) and the coefficient of variation (right), calculated pixel by pixel. The pixel resolution is $0.015\ \arcsec/\text{pixel}$. The axes show the number of pixels, corresponding to 3″across. The colorbar shows the value of each pixel.
COSMOS-Web (Casey et al., 2023, PIs: Kartaltepe & Casey, ID=1727) is a 255-hour JWST treasury imaging survey observing the central area of the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field which is one of the most popular deep multi-wavelength survey fields. It covers a contiguous 0.54 deg2 region in four Near Infrared Camera (NIRCam) filters (F115W, F150W, F277W, and F [^4] W), reaching a $5\sigma$ point source depth of $27.5-28.2$ magnitudes. In parallel, a 0.19 deg2 area of Mid-Infrared Instrument (MIRI) imaging with the F770W filter is covered. For this work, we make use of the 0.28 deg2 JWST/NIRCam F150W images reduced by Zhuang et al. (2024). They used version 1.10.2 of the JWST [^1] https://jwst-pipeline.readthedocs.io/en/latest/ pipeline with the Calibration Reference Data System (CRDS) version of 11.17.0. to reduce the uncalibrated NIRCam raw data retrieved from MAST [^2] https://archive.stsci.edu/doi/resolve/resolve.html?doi=10.17909/6btv-br09. The steps followed in the data reduction process are briefly summarised below:
@@
For each NIRCam mosaic, Zhuang et al. (2024) constructed three different PSF models. For this work, we used the global PSF models, which are produced using all of the point-like sources across the entire field-of-view (FoV) of each dither-combined mosaic. The median PSF full width at half maximum (FWHM) in the COSMOS-Web NIRCam mosaics is 61.1 mas for the F150W filter. The PSF in the NIRCam imaging has a fractional root mean square (RMS) temporal variation of PSF FWHM of $\sim 2.4\%$ for F150W, which is dominated by short timescale fluctuation. In comparison, the spatial variation of PSF FWHM, dominated by random variations, is much larger, at a level of $\geq 5\%$ for short-wavelength filters including F150W. For a full description of their data reduction method, we refer the reader to Sect. 2.1 of Zhuang et al. (2024). The reduced NIRCam F150W imaging data as well as the PSF models used in this work are available at https://ariel.astro.illinois.edu/cosmos_web/. In Fig. 1 we illustrate the variations of the adopted global PSF models. We first stacked all 80 PSF models and then calculated the mean and the standard deviation pixel by pixel (displayed in the top and central panels of Fig. 1, respectively). The bottom panel of Fig. 1 shows the relative dispersion (or coefficient of variation), which is the standard deviation divided by the mean (pixel by pixel).
We use the COSMOS2020 (Weaver et al., 2022) photometric catalogue to construct our sample of real JWST galaxies, within the COSMOS-Web area. For that, we use the Farmer’s version for the photometric catalogue. We use spectroscopic redshifts when available. Otherwise, we adopt the photometric redshifts (photo-$z$) listed in the COSMOS2020 catalogue, computed using the LePhare code (Arnouts et al., 2002; Ilbert et al., 2006). The photo-$z$ precision, given by the normalised median absolute deviation, is around $0.01\times(1+z)$ at $i<24.0$ mag, and $0.03\times(1+z)$ at $27.024.0<i<$ 27.0 mag. Stellar masses ($M_{*}$) are derived by La Marca et al. 2025 (in prep.), using the spectral energy distribution fitting tool CIGALE (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019), including AGN models. These stellar masses are consistent with the stellar masses presented in the COSMOS2020 photometric catalogue obtained by LePhare. We calculate the median bias, $b$, and the median absolute deviation (MAD) between the two stellar mass estimates as follows:
@@
where $\Delta M_{*}=log_{10}M_{*,LePhare}-log_{10}M_{*,CIGALE}$. We find $b=-0.09$ and $MAD=0.013$. We construct our galaxy sample in the redshift range $30.5<z<$ 3, avoiding stars or areas affected by bright stars. Finally, we select a stellar mass complete sample, by including galaxies with $M_{*}$ greater than either $10^{9}\,M_{\odot}$ or the KS-based completeness limit from Weaver et al. (2022) for the COSMOS2020 sample, whichever is higher (the minimum stellar mass limit of $10^{9}\,M_{\odot}$ is imposed to match the limit in the simulated galaxies). In total, we include 25 596 galaxies in our stellar mass complete sample over the 0.28 deg2 analysed in this study.
@@ 2.2 AGN selections
    X-ray AGN selection: For this selection, we use the X-ray photometry data from the Chandra COSMOS Legacy survey (Civano et al., 2016; Marchesi et al., 2016). From the catalogue of optical and infrared counterparts provided by Marchesi et al. (2016), we select sources with a final counterpart identification flag of 1 (secure) or 10 (ambiguous). We then cross-matched this catalogue using the optical coordinates with our selected sources in COSMOS2020. We selected secure X-ray AGN requiring det_ml (the maximum likelihood detection) ¿10.8 in the hard or soft band.
@@
    MIR AGN selection: For this selection, we use the Spitzer/MIPS $24\mu m$ data, provided by the COSMOS-Spitzer programme (Sanders et al., 2007), and the MIR data from the four channels of the Spitzer/IRAC Cosmic Dawn Survey (Euclid Collaboration et al. 2022). We follow Chang et al. (2017) to identify AGN by their MIR emission. We limited the selection to those galaxies with Spitzer/MIPS 24 $\mu$m flux $F_{24\mu m}>20\,\mu$Jy, which is the $1\sigma$ total noise (instrument and confusion noise). Additionally, we required the signal-to-noise ratio ($S/N>5$) in each IRAC channel. Then, we applied the colour--colour criteria in Chang et al. (2017) to select AGN:
@@
    where $x=m_{3.6\mu m}-m_{5.8\mu m}$ and $y=m_{4.5\mu m}-m_{8\mu m}$. Magnitudes are expressed as AB magnitude.
@@ 2.3 Mock JWST images
  • First, we created a smoothed 2D projected map (Rodriguez-Gomez et al., 2019; Martin et al., 2022) constructed from the contribution from all stars after passing them through the JWST/NIRCam F150W filter. These maps do not include a full radiative transfer treatment and therefore do not account for dust. The images were produced with the same pixel resolution as the real JWST observations (0.03 ″/pixel), and have a physical size of $50\times 50$ kpc.
@@
To have a uniform dataset to train our DL model, and to speed up computation, we cut all images to $128\times 128$ pixels, without changing the pixel scale of 0.03″/pixel. In other words, the images are 3.84″across, corresponding to roughly 23 - 33 kpc in physical scale over the redshift range of our sample (i.e. $30.5<z<$ 3).

### v3

@@ Abstract
Methods. We trained a deep learning (DL) based method Zoobot to estimate the fractional contribution of a central point source to the total light. Our training sample comprises realistic mock JWST images of simulated galaxies from the IllustrisTNG cosmological hydrodynamical simulations. We injected different amounts of the observed JWST point spread functions to represent galaxies with varying levels of AGN contribution. Galaxies in our training sample span a wide range of morphologies, including mergers. We analyse in detail the performance of our method as a function of various galaxy properties and compare with results obtained from the traditional light profile fitting tool GALFIT. After training, we applied our method to real JWST observations, in the COSMOS field.
Results. We find excellent performance of our DL method in recovering the injected $f_{AGN}$, in terms of precision and accuracy. The mean difference between the predicted and true $f_{AGN}$ is 0.002-0.002 and the overall root mean square error (RMSE) is $0.013$. The overall relative absolute error (RAE) is $0.076$ and the outlier (defined as predictions with RAE $20\%$) fraction is $6.5\%$. In comparison, using GALFIT, we achieve a mean difference of -$0.02$, RMSE of $0.12$, RAE of $0.19$ and outlier fraction of $19\%$. We also investigate how these key performance metrics obtained from Zoobot and GALFIT vary as a function of the injected $f_{AGN}$, redshift, signal-to-noise ratio, and galaxy size. In addition to the superior performance, our DL method has several other advantages over traditional methods. For example, it has a much higher success rate (even for highly disturbed or irregular galaxies) and is extremely fast. We applied our trained DL model to real JWST observations and found that 20% of the X-ray-selected AGN and 8% of the MIR-selected AGN are also identified as AGN using a cut at $f_{\rm AGN}>0.2$. When using $f_{\rm AGN}>0.1$, these overlaps increase to 33% for the X-ray AGN and 15% for the MIR AGN. In summary, our DL-based method to identify AGN and estimate AGN contribution fraction has a huge potential in future applications to large galaxy imaging surveys.
@@ 1 Introduction
It is crucial to correctly separate the central AGN light from the host galaxy, across a wide range of galaxy types and redshifts. Good quality optical imaging, with high spatial resolution and signal-to-noise ratio (S/N), is needed to decompose the observed total light into contributions from the host galaxy and the central AGN. Traditionally, this is done in photometric data by fitting two-dimensional (2D) profiles to the galaxy’s light, using one (or more if needed) analytic profile to describe the galaxy (typically a Sérsic profile) and a point spread function (PSF) profile to describe the central point source. Many studies use GALFIT (Peng et al., 2002), one of the most widely used 2D surface brightness modelling software to perform image-based decomposition of AGN and host galaxy light (e.g. Kim et al., 2008; Bentz et al., 2009; Gabor et al., 2009; Böhm et al., 2013; Schramm & Silverman, 2013; Du et al., 2014; Urbano-Mayorgas et al., 2019; Son et al., 2022; Aird et al., 2022; Ji et al., 2022; Dewsnap et al., 2023; Zhuang & Ho, 2023; Sturm & Reines, 2024; Zhuang et al., 2024). However, there are some known technical issues with GALFIT. For example, in some cases, the minimisation algorithm used can be trapped in a local minimum, leading to unreliable fits. Other software have been used in an attempt to improve on GALFIT performance, such as PSFMC (Mechtley, 2014), which is a Markov Chain Monte Carlo (MCMC) simultaneous fitting software to perform multi-component profile fitting (Mechtley et al., 2016; Marshall et al., 2021) and LENSTRONOMY (Birrer et al., 2015; Birrer & Amara, 2018), which uses particle swarm optimisation (Kennedy & Eberhart, 1995) for $\chi^{2}$ minimisation, to reduce the likelihood of getting trapped in a local minimum when searching the parameter space, and MCMC for Bayesian parameter inference (Foreman-Mackey et al., 2013). The latter method is used by (Li et al., 2021) to decompose a sample of X-ray-selected AGN into quasars and host galaxy components, to investigate the properties of the host galaxies. However, these approaches all assume that the galaxy’s surface brightness profile can be well fitted by a single Sérsic profile (or combination of Sérsic profiles), which may not be the case, particularly for irregular galaxies, highly disturbed merging galaxies, or galaxies with complicated substructures.
@@
Consequently, if Sérsic profiles cannot adequately describe the host galaxy light, it can introduce a systematic bias in the derived luminosity of the PSF component and of the host galaxy. In addition, if there is a significant contribution to the total flux from the central AGN, it can lead to a bias in the morphology of the host galaxy, as it may appear more bulge-dominated (Pierce et al., 2010). Indeed, it is not always a good idea to represent a galaxy with a single Sérsic profile. Bentz et al. (2009) study AGN host galaxies at redshift $z\approx 0.7$, by fitting their surface brightness distributions with a combination of Sérsic profiles (to account for the host galaxy’s light) and PSF profile (to describe the AGN component). They find that most galaxies in their sample are well described with two Sérsic profiles, in addition to the PSF profile, and a few of them required three or more Sérsic profiles to describe additional components such as bars. Another possible difficulty in decomposing the AGN from the host galaxy is the (sometimes significant) variation of the PSF in a given survey due to spatial/temporal changes or differences in galaxy spectral energy distributions (SEDs), as an incorrect PSF can bias the estimated contribution of the AGN to the total flux of the galaxy. Kim et al. (2008) perform 2D decomposition of Sérsic + PSF profiles to galaxies hosting AGN in Hubble Space Telescope (HST) images and investigate the effect of realistic PSF mismatch, finding a systematic overestimation of the flux of the host galaxies, particularly for those containing bright AGN.
@@
[…]
@@ 2.1 JWST/COSMOS-WEB
Figure 1: Overview of the JWST/NIRCam F150W PSFs in COSMOS-Web. We stacked all available PSFs from Zhuang et al. (2024) and show the mean PSF (left), standard deviation (centre) and the coefficient of variation (right), calculated pixel by pixel. The PSFs have been rebinned to a pixel resolution of $0.03\ \arcsec/\text{pixel}$, matching the resolution used for the synthetic image creation. The axes show the number of pixels, corresponding to 3″across. The colorbar shows the value of each pixel.
COSMOS-Web (Casey et al., 2023, PIs: Kartaltepe & Casey, ID=1727) is a 255-hour JWST treasury imaging survey observing the central area of the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field which is one of the most popular deep multi-wavelength survey fields. It covers a contiguous 0.54 deg2 region in four Near Infrared Camera (NIRCam) filters (F115W, F150W, F277W, and F [^4] W), reaching a $5\sigma$ point source depth of 27.528.227.5-28.2 magnitudes. In parallel, a 0.19 deg2 area of Mid-Infrared Instrument (MIRI) imaging with the F770W filter is covered. For this work, we make use of the 0.28 deg2 JWST/NIRCam F150W images reduced by Zhuang et al. (2024). They used version 1.10.2 of the JWST [^1] https://jwst-pipeline.readthedocs.io/en/latest/ pipeline with the Calibration Reference Data System (CRDS) version of 11.17.0. to reduce the uncalibrated NIRCam raw data retrieved from MAST [^2] https://archive.stsci.edu/doi/resolve/resolve.html?doi=10.17909/6btv-br09. The steps followed in the data reduction process are briefly summarised below:
@@
For each NIRCam mosaic, Zhuang et al. (2024) constructed three different PSF models. For this work, we used the global PSF models, which are produced using all of the point-like sources across the entire field-of-view (FoV) of each dither-combined mosaic. The median PSF full width at half maximum (FWHM) in the COSMOS-Web NIRCam mosaics is 61.1 mas for the F150W filter. The PSF in the NIRCam imaging has a fractional root mean square (RMS) temporal variation of PSF FWHM of $\sim 2.4\%$ for F150W, which is dominated by short timescale fluctuation. In comparison, the spatial variation of PSF FWHM, dominated by random variations, is much larger, at a level of $\geq 5\%$ for short-wavelength filters including F150W. For a full description of their data reduction method, we refer the reader to Sect. 2.1 of Zhuang et al. (2024). The reduced NIRCam F150W imaging data as well as the PSF models used in this work are available at https://ariel.astro.illinois.edu/cosmos_web/. In Fig. 1 we illustrate the variations of the adopted global PSF models. We first stacked all 80 PSF models and then calculated the mean and the standard deviation pixel by pixel (displayed in the left and central panels of Fig. 1, respectively). The right panel of Fig. 1 shows the relative dispersion (or coefficient of variation), which is the standard deviation divided by the mean (pixel by pixel).
We use the COSMOS2020 (Weaver et al., 2022) photometric catalogue to construct our sample of real JWST galaxies, within the COSMOS-Web area. For that, we use the Farmer’s version for the photometric catalogue. We use spectroscopic redshifts when available. Otherwise, we adopt the photometric redshifts (photo-$z$) listed in the COSMOS2020 catalogue, computed using the LePhare code (Arnouts et al., 2002; Ilbert et al., 2006). The photo-$z$ precision, given by the normalised median absolute deviation, is around $0.01\times(1+z)$ at $i<24.0$ mag, and $0.03\times(1+z)$ at $24.0<i<27.0$ mag. Stellar masses ($M_{*}$) are derived by La Marca et al. 2025 (in prep.), using the spectral energy distribution fitting tool CIGALE (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019), including AGN models. These stellar masses are consistent with the stellar masses presented in the COSMOS2020 photometric catalogue obtained by LePhare. We calculate the median bias, $b$, and the median absolute deviation (MAD) between the two stellar mass estimates as follows:
@@
where $\Delta M_{*}=log_{10}M_{*,LePhare}-log_{10}M_{*,CIGALE}$. We find b=0.09b=-0.09 and $MAD=0.013$. We construct our galaxy sample in the redshift range $0.5<z<3$, avoiding stars or areas affected by bright stars. Finally, we select a stellar mass-complete sample by including galaxies with $M_{*}$ greater than either $10^{9}\,M_{\odot}$ or the $K_{S}$-based completeness limit from Weaver et al. (2022) for the COSMOS2020 catalog:
$M_{\mathrm{lim}}(z)$ $=-3.55\times 10^{8}(1+z)$
$\quad+2.70\times 10^{8}(1+z)^{2},$ (3)
whichever is higher. The minimum stellar mass limit of $10^{9}\,M_{\odot}$ is imposed to match the limit used for the simulated galaxy sample. In total, we include 25 596 galaxies in our stellar mass complete sample over the 0.28 deg2 analysed in this study.
@@ 2.2 AGN selections
    X-ray AGN selection: For this selection, we use the X-ray photometry data from the Chandra COSMOS Legacy survey (Civano et al., 2016; Marchesi et al., 2016). From the catalogue of optical and infrared counterparts provided by Marchesi et al. (2016), we select sources with a final counterpart identification flag of 1 (secure) or 10 (ambiguous). We then cross-matched this catalogue using the optical coordinates with our selected sources in COSMOS2020. We selected secure X-ray AGN requiring det_ml (the maximum likelihood detection) ¿10.8 in the hard or soft band. This selection yields a total of 223 X-ray AGN within our final sample.
@@
[…]
@@
    where $x=m_{3.6\mu m}-m_{5.8\mu m}$ and $y=m_{4.5\mu m}-m_{8\mu m}$. Magnitudes are expressed as AB magnitude. Applying these criteria results in 680 MIR-selected AGN in our final sample.
@@ 2.3 Mock JWST images
  • First, we created a smoothed 2D projected map (Rodriguez-Gomez et al., 2019; Martin et al., 2022). For each stellar particle in the simulation, we assigned a SED based on its mass, age, and metallicity, using the (Bruzual & Charlot, 2003) stellar population synthesis models, with a Chabrier (2003) initial mass function. These SEDs were integrated through the JWST/NIRCam F150W filter transmission curve to obtain the flux contribution of each particle in that band. The spatial distribution of flux was smoothed using an adaptive kernel to create the 2D map. This approach does not include full radiative transfer calculations, and therefore dust attenuation and emission lines are not modeled. The images were produced with the same pixel resolution as the real JWST observations (0.03 ″/pixel), and have a physical size of $50\times 50$ kpc.
@@
To have a uniform dataset to train our DL model, and to speed up computation, we cut all images to $128\times 128$ pixels, without changing the pixel scale of 0.03″/pixel. In other words, the images are 3.84″across, corresponding to roughly 23 - 33 kpc in physical scale over the redshift range of our sample (i.e. $0.5<z<3$).